Amalgamated General Corporation is a consulting firm that also offers financial services through its credit division. From time to time the company buys and sells securities. The following selected transactions relate to Amalgamated’s investment activities during the last quarter of 2018 and the first month of 2019. The only securities held by Amalgamated at October 1 were $50 million of 10% bonds of Kansas Abstractors, Inc., purchased on May 1 at face value and held in Amalgamated’s trading portfolio. The company’s fiscal year ends on December 31.

2018
Oct. 18 Purchased 2 million preferred shares of Millwork Ventures Company for $74 million.
31 Received semiannual interest of $2.0 million from the Kansas Abstractors bonds.
Nov. 1 Purchased 10% bonds of Holistic Entertainment Enterprises at their $72 million face value, to be held until they mature in 2025. Semiannual interest is payable April 30 and October 31.
1 Sold the Kansas Abstractors bonds for $45 million because rising interest rates are expected to cause their fair value to continue to fall. No unrealized gains and losses had been recorded on these bonds previously.
Dec. 1 Purchased 12% bonds of Household Plastics Corporation at their $80 million face value, to be held until they mature in 2028. Semiannual interest is payable May 31 and November 30.
20 Purchased U. S. Treasury bonds for $6.7 million as trading securities, hoping to earn profits on short-term differences in prices.
21 Purchased 4 million common shares of NXS Corporation for $66 million, planning to earn profits from dividends or gains if prevailing market conditions encourage sale.
23 Sold the Treasury bonds for $7.1 million.
29 Received cash dividends of $3 million from the Millwork Ventures Company preferred shares.
31 Recorded any necessary adjusting entry(s) and closing entries relating to the investments. The market price of the Millwork Ventures Company preferred stock was $36.00 per share and $19.00 per share for the NXS Corporation common. The fair values of the bond investments were $54.6 million for Household Plastics Corporation and $17.8 million for Holistic Entertainment Enterprises.
2019
Jan. 7 Sold the NXS Corporation common shares for $61 million.
Answer #1

Ans)

Journal Entries
for the year 2018
Date Particulers Debit (millions) Credit (millions)
2018
18th Oct Invest in preferred shares of Millwork Ventures Company a/c 74
   To Cash a/c 74
(being purchased preferred share)
31st Oct Cash a/c 2
   To Investment Revenue a/c(c) 2
(Being Received semiannual interest)
1st Nov Investment in Holistic Entertainment Enterprises a/c 72
   To Cash a/c 72
(Being Purchased 10% bonds of Holistic Entertainment Enterprises )
1st Nov Cash a/c 45
Loss on Sale a/c 5
Investment in Kansas Abstractors bonds a/c 50
(being Sold the Kansas Abstractors bonds)
1st Dec Investment in Household Plastics Corporation a/c 80
To cash a/c 80
(being Purchased 12% bonds of Household Plastics Corporation )
20th Dec Investment in U S Treasury bonds a/c 6.7
To cash a/c 6.7
(Being Purchased U. S. Treasury bonds )
21st Dec Investment in NXS Corporation a/c 66
To Cash a/c 66
(being Purchased 4 million common shares of NXS Corporation)
23rd Dec Cash a/c 7.1
To Investment in US Treasury Bonds a/c 6.7
To Gain on Sale a/c 0.4
(being sold the US treasury bonds in profit)
29th Dec Cash a/c 3
To Investment Revenue a/c(a) 3
(Being Received cash dividends )
31st Jan Investment Revenue Receivable from Household Plastics Corporation a/c (i) 0.8
Investment revenue Receivable from Holistic Entertainment Enterprise a/c(ii) 1.2
To Investment Revenue a/c(b) 2
(being investment revenue receivable)

Calculations;

(i) Revenue Receivable from Household Plastics Corporation; 12% * 80 * 1month= 12/100 * 80 * 1/12 = 0.8

(i) Revenue Receivable from Holistic Entertainment Enterprise; 10% * 72 * 2months= 10/100 * 72 * 2/12= 1.2

Adjusting Entries-

31st Dec Net holding Unrealised Gains/Losses a/c (i) 2
To Fair Value adjustment a/c 2
(Being value adjusted)
31st Dec Fair Value Adjustment a/c 5
To Net Holding unrealised Gains/Losses a/c 5
(Being Value adjusted )

Closing Entries-

31st Jan Net holding Unrealised Gains/Losses a/c 5
Investment Revenue a/c (a+b+c)=(3+2+2) 7
Gain on Sale of Investment a/c 0.4
To Loss on sale of Investment a/c 5
To Income Summary a/c 7.4
2019
7th Jan Cash a/c 61
loss on Sale a/c 5
To Investment in NXS a/c 66